OpenStudy (anonymous):

what number equals (sqrt)1,377

OpenStudy (lgbasallote):

this is better asked to a scientific calculator....:)

OpenStudy (lgbasallote):

but just for fun....37.10795063 i did that mentally ;D

OpenStudy (mertsj):

\[\sqrt{1377}=\sqrt{81(17)}=9\sqrt{17}\]

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):